okay, i just dont know how to graph this problem.
|n+8|>=3
my work was:
n+8>=3 n+8>=-3
-8 -8 -8 -8
n>=-5 n>=-11
now what?

Answer by ankor@dixie-net.com(22740)

|n+8| > = 3
my work was:
n + 8 > = 3; this right, so you have;
n >= 3 -8
n >= -5 which you have
and
n + 8 <= -3; when you make 3 negative, you have to reverse the inequality sign
n <= -3 - 8
n <= -11
:
The solution:
n >= -5
n <= -11
:
:
Note that any value between -11 and -5 would not satisfy the inequality
But any value below -11 and above -5 would satisfy